OpenStudy (anonymous):

(edit: I sill need help, despite length of post)Can anybody check my answers? (geometry/trig)

OpenStudy (anonymous):

Which pair of angles has congruent values for the sin x° and the cos y°? 70°; 160° 70°; 70 70°; 120° 70°; 20°

OpenStudy (anonymous):

I said D.

OpenStudy (anonymous):

Calculate the area of triangle ABC with altitude CD, given A (−3, −4), B (−6, 2), C (0, 0), and D (−4, −2). 14 square units 15 square units 18 square units 20 square units

OpenStudy (anonymous):

I said A.

OpenStudy (campbell_st):

well there is a property is triangles perhaps this diagram may help 

OpenStudy (anonymous):

sin(x)= o/h ; cos(90-x)=a/h is that what you mean??

OpenStudy (campbell_st):

thats it... cos(x) = sin(90 - x)

OpenStudy (anonymous):

so D would be right?

OpenStudy (anonymous):

@amistre64 @e.mccormick @wolfe8 ?

Parth (parthkohli):

Apparently not. Do you see two angles that have a difference of 90 degrees?

OpenStudy (anonymous):

oh. for some reason I was thinking it should = 90 degrees

OpenStudy (anonymous):

so it'd be A, but was the image they added relevant?

Parth (parthkohli):

Yes.

Parth (parthkohli):

It showed you how cos(90 - x) = sin(x).

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

thanks. what about my second question?